Fermi National Laboratory
 

Accelerator Update for May 30, 2007 - June 1, 2007

 

Wednesday May 30

The day shift began with Operations monitoring Tevatron (TeV) store 5467 (48.08E30) and the Recycler stash (213.1E10), with the Antiproton Source (Pbar) stacking (29.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

The Neutron Therapy Facility (NTF) reported at 11:19 AM that they had completed treatment on their last patient for the day.

At 12:12 PM, a Booster power supply (GMPS) tripped off on a ground fault indication. Experts investigated, cleared the ground fault, reset a regulator card, and had Booster sending beam by 1:31 PM.

Pbar experts began a study at 2:44 PM that made changes to the Debuncher Accumulator line.

TeV experts began an end-of-store study at 3:06 PM. Another expert began a Tevatron Electron Lens (TEL1) study.

The Pbar study ended at 3:39 PM.

Operations reported at 4:18 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations terminated store 5467 at 5:03 PM.

Operations put the TeV into shot setup at 5:18 PM.

At 6:22 PM, all the MI VCBs opened. Operators investigated and discovered that a ground fault on the MI-60N quadrupole bus had caused the MI power supply interlock loop to trip. Operators had everything reset and MI back on in about 15 minutes.

At 7:44 PM, the Recycler suffered a vacuum burst that spread. A Recycler expert investigated. The Recycler lost 8E10 antiprotons during the burst. It recovered.

Operations established store 5468 at 7:51 PM with an initial luminosity of 238.2E30.

At 10:43 PM, a Linac Klystron RF station (KRF5) tripped off for the first of three times in about an hour. It reset.

At 11:53 PM, the 750keV line buncher tripped off. It reset.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Thursday May 31

The midnight shift began with Operations monitoring store 5468 (129.37E30) and the stash (48.5E10), with Pbar stacking (51.3E10), and with NuMI and MiniBooNE taking beam.

Operations reported at 2:04 AM that the normal Recycler stochastic cooling was going too slow. Operators turned on the Pelletron.

At 6:30 AM, Operations reported that earlier in the shift TEL1 tripped off on a bad vacuum indication. Operators reset the vacuum interlock chassis and the QPM logic module and TEL1 recovered.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

A MI expert began a slipstacking study at 10:13 AM. She had completed her study by 12:11 PM.

A MI expert began a barrier bucket study at 1:26 PM.

At 3:29 PM, Operations reported that TEL2 tripped off due to bad vacuum and quench bit indications. Operators reset it, but it tripped again. Operations contacted experts.

At 4 PM, Operations halted the MiniBooNE beam due to a problem with their Data Acquisition system (DAQ). MiniBooNE experts investigated.

Operations reported at 4:06 PM that the Recycler had stashed three times during the day shift.

Operations reported at 4:26 PM that TEL2 was back on. Experts reported that a TeV corrector power supply (T:L2C3) tripped and caused a vacuum trip that in turn caused TEL2 to trip. They also thought that there might have been a QPM problem.

At 5:32 PM, TEL2 tripped off. It reset.

Operations reported at 5:34 PM that the MiniBooNE DAQ problem was resolved. MiniBooNE resumed taking beam.

The MI expert completed her barrier bucket study at 5:53 PM.

Operations reported at 11:58 PM that the Recycler had stashed three times during the evening shift.

Friday June 1

The midnight shift began with Operations monitoring store 5468 and the stash (337.4E10), with Pbar stacking (24.2E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 5468 at 12:04 AM.

Operations put the TeV into shot setup at 12:23 AM.

At 12:27 AM, the NuMI beam permit tripped off due to beam in the kicker gap. It reset.

Operations established store 5469 at 2:06 AM with an initial luminosity of 237.59E30.

Operations reported at 8 AM that the Recycler had stashed two times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. However, the Run Coordinator will allow Linac to conduct a LRF3 tune that will disrupt the beam. During this tune, Booster will work on GMPS. Both Pbar and MI experts will be allowed to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/01/2007    email Fermilab